The design of a concrete mix requires 1173 kg/m3



of gravel in dry condition, 582 kg/m³ of sand in dry condition, and 157 kg/m³of free water. The gravel available at the job site has a moisture content of 0.8% and absorption of 1.5%, and the available sand has a moisture content of 1.1% and absorption of 1.3%. What are the masses of gravel, sand, and water per cubic meter that should be used at the job site?

1
Expert's answer
2022-03-17T14:55:45-0400

Answer:


- the weight of the gravel is 1183.56 kg/m³


- the weight of sand is 588.98 kg/m³


- the weight of water is 165.784 kg/m³


Explanation:


 Given the data in the question;


mixture content of the gravel = 1173 kg/m³


sand content = 582 kg/m³


free water content = 157 kg/m³


moist of gravel = 0.9%


absorption of gravel = 1.5%


moist of sand = 1.2%


absorption of sand = 1.5% 


masses of gravel, sand, and water per cubic meter that should be used at the job site ?




To get the weight of moist gravel, we use the following equation;


Weight of moist gravel = content of gravel × ( 1 + moist of gravel )


we substitute


Weight of moist gravel = 1173 × ( 1 + 0.9%)


Weight of moist gravel = 1173 × ( 1 + 0.009)


Weight of moist gravel = 1173 × 1.009


Weight of moist gravel = 1183.56 kg/m³


Therefore, the weight of the gravel is 1183.56 kg/m³


To get weight of the sand, we use the following formula;


Weight of sand = sand content × ( 1 + moist of sand )


we substitute


Weight of sand = 582 × ( 1 + 1.2% )


Weight of sand = 582 × ( 1 + 0.012 )


Weight of sand = 582 × 1.012


Weight of sand = 588.98 kg/m³


Therefore, the weight of sand is 588.98 kg/m³



Weight of mix water= "mix-Water weight"


= 157-1173(0.9%-1.5%) - 582(1.2% - 1.5%)


= 157+7.038+1.746


= 165.784kg/m3


Therefore, the weight of water is 165.784kg/m3